OpenStudy (anonymous):

pi/2 + 4 = in radians

OpenStudy (anonymous):

\[\pi/2 +4 \] radians. What answer are you looking for?

OpenStudy (anonymous):

Im working on a section which deals with trigonometric graphs, amplitude, period, etc etc and it also has shifts. For this certain problem i have to add 4 to pi/2 then plot it on a coordinate plane.

OpenStudy (anonymous):

WHOA! ARE YOU IN MY CLASS? I ASKED THIS EARLIER!

OpenStudy (anonymous):

You still haven't clarified your question. An angle in radians is just a number, 4+pi/2 is just 4+pi/2. You can't really simplify it much. There must be some context to the question you have not included.

OpenStudy (anonymous):

Ok how about i just show you the original problem..... y=sin 3x-4

OpenStudy (anonymous):

\[(\pi+8)/2\]

OpenStudy (anonymous):

What are you solving for?

OpenStudy (anonymous):

"Graph the sine or cosine function"

OpenStudy (anonymous):

Ahaa. In which case, you will be drawing a sine wave, which crosses the y-axis at y=-4, and has a period of 3 cycles per 2pi. See here: http://www.wolframalpha.com/input/?i=y%3Dsin+3x-4+

OpenStudy (anonymous):

Could you thoroughly explain the steps, just looking at the graph does not help me at all.

OpenStudy (anonymous):

So, when you look at the graph y=sin(x), how long does it take to get through one whole wave (positive and negative parts), starting at x=0?

OpenStudy (anonymous):

I'm not sure what you're trying to say but my teacher said to first find a = amplitude, b, period = 2pi/b and make a x|y chart to determine the coordinates. Then, add the x and y by whatever the shift is (h,k). I can't graph this without knowing what the pattern for the x coordinate is which is why i asked what pi/2 + 4 is.

OpenStudy (dumbcow):

i think you might mean sin(3x-4) http://www.wolframalpha.com/input/?i=y%3Dsin+%283x-4%29+ as far as pi/2 + 4, just estimate with calculator

OpenStudy (anonymous):

If indeed that is what you mean't, it would indeed be a rather different question.

OpenStudy (anonymous):

He wants it in radians not decimals and the book just says y=sin 3x -4 so sorry for any confusion.

OpenStudy (anonymous):

Ok, so the 3 determines your period (the wave will have a three times higher frequency than the normal sine wave) and the four will be a phase shift (imagine picking up the wave, and moving it 4 to the right). Because of that 4, you won't get any clean radian measures. I think you may be confused, degrees are not the same thing as decimals. Radians may be expressed as decimals. He probably just wants to make certain you are using radians in stead of degrees. For that, you really should just use a calculator set in radian mode.

OpenStudy (dumbcow):

amplitude = 1 period = 2pi/3 horizontal shift = 4/3 -> 3x-4 = 0 x = 4/3

OpenStudy (anonymous):

@dumbcow: I understood until your horizontal shift, there is no horizontal shift. It's a vertical shift and it's 4. y=asinb(x-h)+k

OpenStudy (dumbcow):

it depends on if its sin(3x)-4 or sin(3x-4) sin(3x-4) = sin(3(x-4/3)) so h = 4/3

OpenStudy (anonymous):

As i said, the book has it exactly written as y= sin 3x-4

OpenStudy (anonymous):

@dfyodor: Im getting so confused, so just graph it normally and then move it to the right 4 units?